how much heat is given out when 50g of steam at 100°C cool to water at 28°C?

1
Expert's answer
2022-03-03T14:24:02-0500

Answer

"Heat is given out\\\\ Q = m\\> C_{p}\\times \\Delta t\\\\Q=m\u00d7C \np\n\u200b\t\n \u00d7\u0394t\n\n\n\nQ = 0.05 \\times 2.3\\times10^{-3}\\times (1000-280) =82800 kJQ=0.05\u00d72.3\u00d710 \n\u22123\n \u00d7(1000\u2212280)\\\\=82800kJ"